OpenStudy (anonymous):

Evaluate: f(x)=2sinx+4cos2x @x=π/6

OpenStudy (anonymous):

x = π/6 f(π/6) = 2sin(π/6) + 4cos(π/3) f(π/6) = 2 * 1/2 + 4 * 1/2 f(π/6) = 1 + 2 = 3